A researcher believes that 8%

of pet dogs in Europe are Labradors.

If the researcher is right, what is the probability that the proportion of Labradors in a sample of 437

pet dogs would differ from the population proportion by greater than 3%? Round your answer to four decimal places.
Answer #1

P(Sample proportion would differ by more than 0.03)

= 1 - P(Within 0.03)

\small = 1 - P(-\frac{0.03}{\sqrt{0.08*0.92/437}} < z < \frac{0.03}{\sqrt{0.08*0.92/437}})

= 1 - P(-2.31 < z < 2.31)

= 1 - 0.9792

= 0.0208